2. A child pulls a wagon by the handle along a flat sidewalk. She exerts a force of 76.0 N at an angle of 27 degrees ° above the horizontal while she moves the wagon 22.0 m forward. The force of friction on the wagon is 17.0 N. a) Calculate the mechanical work done by the child on the wagon. b) Calculate the total work done on the wagon.


1
Expert's answer
2021-01-10T18:28:36-0500

a) We can calculate the mechanical work done by the child on the wagon from the formula:


"W=Fdcos\theta=76.0\ N\cdot 22.0\ m\cdot cos27^{\circ}=1490\ J."

b) The total work done on the wagon can be found as follows:


"W_{tot}=W+W_{fr},""W_{tot}=1490 J+(-17.0\ N)\cdot 22.0\ m=1116\ J."

Answer:

a) "W=1490\ J."

b) "W_{tot}=1116\ J."